# Exclusive $J/\psi$, $\psi(2s)$, and $e^{+}e^{-}$ pair production in Au$+$Au ultra-peripheral collisions at RHIC

### v1

@@ VI.1 Incoherent interactions and nuclear breakup
Figure 10 shows the differential cross section of $J/\psi~{}$photoproduction as a function of $p^{2}_{\mathrm{T}}~{}$in the full rapidity range $|y|<1$ for the $all~{}n$ case; the subset $0n0n$ neutron category is also shown. In the coherent peak at lowest $p^{2}_{\mathrm{T}}$, the ratio of $0n0n$ to $all~{}n$ is $\sim 40\%$, consistent with the fraction of $0n0n$ photon flux, where no neutrons are produced by nuclear dissociation following Coulomb excitation. At higher $p^{2}_{\mathrm{T}}$, where the incoherent processes dominates, the additional mechanisms for neutron emission result in the $0n0n$ fraction dropping to 10-20%.
@@
This is confirmed by the highly asymmetric rapidity distribution for incoherent $J/\psi~{}$photoproduction as shown in Fig. 11. The figure also shows the result from a BeAGLE modelling of this reaction. The model provides a good description of the shape of the data. The difference in normalization may be due to the shortcomings of BeAGLE cited earlier. This asymmetric shape was also described in an earlier discussion of the LTA model [17].
@@ VI.5 $\gamma\gamma\rightarrow e^{+}e^{-}$ cross sections
In Fig. 15, the differential cross sections of $\gamma\gamma$ to $e^{+}e^{-}$ pairs, as a function of the pair mass $m_{\rm{ee}}$, are shown for different neutron configurations. The data are compared with both STARlight and a QED calculation performed by Zha et al [38, 37]. The ratios between the data and these two predictions are shown in Fig. 16. The central value of the data are 10-20% above the STARlight prediction, and 10-20% below the QED calculation. It should be noted that STARlight does not include $e^{+}e^{-}$ pair production inside the nucleus, whereas the QED calculation does. The magnitude of this effect has been estimated to be $\sim 10\%$ [2], partially accounting for the discrepancy seen in STARlight relative to the data. The scale uncertainty from the luminosity measurement is 10% (not shown), implying that both models are consistent with the data.
@@
Here $\mathcal{M}$, the QED matrix element for $\gamma\gamma\rightarrow e^{+}e^{-}$, is well known. The $\gamma\gamma$ cross section is quadratic in the photon fluxes, while the Au$+$Au cross sections are linear in the flux. Therefore, the $\sigma_{\gamma\gamma}$ is a test of the photon fluxes. The 10-20% model and data discrepancy in Fig. 16 implies that deviations in $\sigma_{\rm Au+Au}$ are half as large. Therefore, the fluxes used in the $\gamma$+Au cross sections measurement are valid at the 5-10% level.
@@
[…]
@@ VII Conclusion
Exclusive $J/\psi$, $\psi(2s)$, and $e^{+}e^{-}$ pair photoproduction in Au$+$Au UPCs at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV using the STAR detector are measured. For $J/\psi~{}$photoproduction, both coherent and incoherent processes as a function of rapidity $y$, $p^{2}_{\mathrm{T}}$, and different neutron configurations are presented. In particular, three different neutron configurations, $0n0n$, $0nXn$, and $XnXn$, are combined to resolve the photon energy ambiguity in UPCs, which leads to the total cross section of coherent $J/\psi~{}$photoproduction in $\gamma^{\ast}+$Au collisions as a function of photon-nucleon center-of-mass energy. It is found that the coherent nuclear suppression factor at $W_{\mathrm{\gamma*N}}~{}$$=25.0$ GeV is $0.71\pm 0.10$ when compared to the expectation of a free nucleon. This suppression supports the nuclear shadowing effect with the leading twist approximation. The Next-to-Leading order calculation of perturbative Quantum Chromodynamics on coherent $J/\psi~{}$photoproduction is compared with this measurement. The description of the data is off by a factor of two at midrapidity based on nPDF EPPS21, which may indicate the large uncertainty on the nuclear parton distribution functions that this data can significantly constrain. The nuclear parton density at the top RHIC energy is at the region between large momentum quarks ($x_{\rm{parton}}>0.1$) and low momentum gluons ($x_{\rm{parton}}<0.001$), which is essential to the understanding of nuclear modification effects in this transition regime. Moreover, incoherent $J/\psi~{}$photoproduction has been measured up to high $p^{2}_{\mathrm{T}}~{}$of $2.2~{}\rm{(GeV/c)^{2}}$, and the incoherent suppression factor at $W_{\mathrm{\gamma*N}}~{}$$=25.0$ GeV is found to be $0.36\pm 0.07$ relative to the free proton. Based on a hot-spot model with sub-nucleonic parton density fluctuation, the incoherent data indicate a similar level of fluctuation seen in the free proton as it is characterized by the shape of the $p^{2}_{\mathrm{T}}~{}$distribution. However, direct comparisons between hot-spot models (Sartre and CGC) and data cannot be fully reconciled and further theory investigations are needed to draw a conclusion. Finally, the QED $\gamma\gamma\rightarrow e^{+}e^{-}$ has been measured up to an invariant mass of 6 $\rm GeV/c^{2}$ for different neutron emission classes, which constrains the modelling of neutron emission and photon flux. The data provide important constraints to the parton density and its fluctuations, and also provide an essential experimental baseline for such measurement at the upcoming Electron-Ion Collider.
@@ Acknowledgements
We thank the RHIC Operations Group and RCF at BNL, the NERSC Center at LBNL, and the Open Science Grid consortium for providing resources and support. This work was supported in part by the Office of Nuclear Physics within the U.S. DOE Office of Science, the U.S. National Science Foundation, National Natural Science Foundation of China, Chinese Academy of Science, the Ministry of Science and Technology of China and the Chinese Ministry of Education, the Higher Education Sprout Project by Ministry of Education at NCKU, the National Research Foundation of Korea, Czech Science Foundation and Ministry of Education, Youth and Sports of the Czech Republic, Hungarian National Research, Development and Innovation Office, New National Excellency Programme of the Hungarian Ministry of Human Capacities, Department of Atomic Energy and Department of Science and Technology of the Government of India, the National Science Centre and WUT ID-UB of Poland, the Ministry of Science, Education and Sports of the Republic of Croatia, German Bundesministerium für Bildung, Wissenschaft, Forschung and Technologie (BMBF), Helmholtz Association, Ministry of Education, Culture, Sports, Science, and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) and Agencia Nacional de Investigación y Desarrollo (ANID) of Chile.

### v3

@@ VI.1 Incoherent interactions and nuclear breakup
Figure 10 shows the differential cross section of $J/\psi~{}$photoproduction as a function of $p^{2}_{\mathrm{T}}~{}$in the full rapidity range $|y|<1$ for the $all~{}n$ case; the subset $0n0n$ neutron category is also shown. In the coherent peak at lowest $p^{2}_{\mathrm{T}}$, the ratio of $0n0n$ to $all~{}n$ is ≈$\approx 40\%$, consistent with the fraction of $0n0n$ photon flux, where no neutrons are produced by nuclear dissociation following Coulomb excitation. At higher $p^{2}_{\mathrm{T}}$, where the incoherent processes dominates, the additional mechanisms for neutron emission result in the $0n0n$ fraction dropping to 10-20%.
@@
This is confirmed by the highly asymmetric rapidity distribution for incoherent $J/\psi~{}$photoproduction as shown in Fig. 11. The figure also shows the result from a BeAGLE modelling of this reaction. The model provides a good description of the shape of the data. The difference in normalization may be due to the shortcomings of BeAGLE cited earlier. This asymmetric shape was also described in an earlier discussion of the LTA model [17]. Although early LHC data have seen this qualitative behavior [3], the STAR data has shown it explicitly for the first time, especially for the incoherent production.
@@ VI.5 $\gamma\gamma\rightarrow e^{+}e^{-}$ cross sections
In Fig. 15, the differential cross sections of $\gamma\gamma$ to $e^{+}e^{-}$ pairs, as a function of the pair mass $m_{\rm{ee}}$, are shown for different neutron configurations. The data are compared with both STARlight and a QED calculation performed by Zha et al [39, 38]. The ratios between the data and these two predictions are shown in Fig. 16. The central value of the data are 10-20% above the STARlight prediction, and 10-20% below the QED calculation. It should be noted that STARlight does not include $e^{+}e^{-}$ pair production inside the nucleus, whereas the QED calculation does. The magnitude of this effect has been estimated to be ≈$\approx 10\%$ [2], partially accounting for the discrepancy seen in STARlight relative to the data. The scale uncertainty from the luminosity measurement is 10% (not shown), implying that both models are consistent with the data.
@@
Here $\mathcal{M}$, the QED matrix element for $\gamma\gamma\rightarrow e^{+}e^{-}$, is well known. The $\gamma\gamma$ cross section is quadratic in the photon fluxes, while the Au$+$Au cross sections are linear in the flux. Therefore, the $\sigma_{\gamma\gamma}$ is a test of the photon fluxes. The 10-20% model and data discrepancy in Fig. 16 implies that deviations in $\sigma_{\rm Au+Au}$ are half as large. Therefore, the fluxes used in the $\gamma$+Au cross sections measurement are valid at the 5-10% level. Other models, such as SuperChic 3 [64, 65, 66], may be compared with the reported data in the future.
@@
Figure 15 also shows the $\gamma\gamma\rightarrow e^{+}e^{-}$ cross section for the $XnXn$ category from a previous STAR publication [60]. That measurement at lower $m_{\rm{ee}}$ utilized different experimental techniques than the present data, based on TOF as opposed to BEMC selections. The two data sets have excellent agreement in the overlap region near $m_{\rm{ee}}\approx 2$ GeV/c2, providing further affirmation of the measurements. These results present the first $\gamma\gamma\rightarrow e^{+}e^{-}$ measurement up to an invariant mass of 6 $\rm GeV/c^{2}$.
@@ VII Conclusion
Exclusive $J/\psi$, $\psi(2s)$, and $e^{+}e^{-}$ pair photoproduction in Au$+$Au UPCs at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV using the STAR detector are measured. For $J/\psi~{}$photoproduction, both coherent and incoherent processes as a function of rapidity $y$, $p^{2}_{\mathrm{T}}$, and different neutron configurations are presented. In particular, three different neutron configurations, $0n0n$, $0nXn$, and $XnXn$, are combined to resolve the photon energy ambiguity in UPCs, which leads to the total cross section of coherent $J/\psi~{}$photoproduction in $\gamma+$Au collisions as a function of photon-nucleon center-of-mass energy. It is found that the coherent nuclear suppression factor at $W_{\mathrm{\gamma N}}~{}$$=25.0$ GeV is $0.71\pm 0.10$ when compared to the expectation of a free nucleon. This suppression supports the nuclear shadowing effect with the leading twist approximation. The Next-to-Leading order calculation of perturbative Quantum Chromodynamics on coherent $J/\psi~{}$photoproduction is compared with this measurement. The description of the data is off by a factor of two at midrapidity based on nPDF EPPS21, which may indicate the large uncertainty on the nuclear parton distribution functions that this data can significantly constrain. The nuclear parton density at the top RHIC energy is at the region between large momentum quarks ($x_{\rm{parton}}>0.1$) and low momentum gluons ($x_{\rm{parton}}<0.001$), which is essential to the understanding of nuclear modification effects in this transition regime. Moreover, incoherent $J/\psi~{}$photoproduction has been measured up to high $p^{2}_{\mathrm{T}}~{}$of $2.2~{}\rm{[(GeV/c)^{2}]}$, and the incoherent suppression factor at $W_{\mathrm{\gamma N}}~{}$$=25.0$ GeV is found to be $0.36\pm 0.07$ relative to the free proton. Based on a hot-spot model with sub-nucleonic parton density fluctuation, the incoherent data indicate a similar level of fluctuation seen in the free proton as it is characterized by the shape of the $p^{2}_{\mathrm{T}}~{}$distribution. However, direct comparisons between hot-spot models (Sartre and CGC) and data cannot be fully reconciled and further theory investigations are needed to draw a conclusion. Finally, the QED $\gamma\gamma\rightarrow e^{+}e^{-}$ has been measured up to an invariant mass of 6 $\rm GeV/c^{2}$ for different neutron emission classes, which constrains the modelling of neutron emission and photon flux. The data provide important constraints to the parton density and its fluctuations, and also provide an essential experimental baseline for such measurement at the upcoming Electron-Ion Collider.
@@ Acknowledgements
We thank the RHIC Operations Group and RCF at BNL, the NERSC Center at LBNL, and the Open Science Grid consortium for providing resources and support. This work was supported in part by the Office of Nuclear Physics within the U.S. DOE Office of Science, the U.S. National Science Foundation, National Natural Science Foundation of China, Chinese Academy of Science, the Ministry of Science and Technology of China and the Chinese Ministry of Education, the Higher Education Sprout Project by Ministry of Education at NCKU, the National Research Foundation of Korea, Czech Science Foundation and Ministry of Education, Youth and Sports of the Czech Republic, Hungarian National Research, Development and Innovation Office, New National Excellency Programme of the Hungarian Ministry of Human Capacities, Department of Atomic Energy and Department of Science and Technology of the Government of India, the National Science Centre and WUT ID-UB of Poland, the Ministry of Science, Education and Sports of the Republic of Croatia, German Bundesministerium fĂźr Bildung, Wissenschaft, Forschung and Technologie (BMBF), Helmholtz Association, Ministry of Education, Culture, Sports, Science, and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) and Agencia Nacional de InvestigaciĂłn y Desarrollo (ANID) of Chile.